OpenStudy (anonymous):

h(t)=cot t intervals: pie/4, 3pie/4 totally lost on this one... cant find a formula to plug in the intervals. Help :/ Thanks.

OpenStudy (anonymous):

are you supposed to graph?

OpenStudy (anonymous):

\[\cot t = 1/\tan t\] So, cotangent is just an inverse of tangent. Hope this helps.

OpenStudy (anonymous):

not "inverse" but "reciprocal"

OpenStudy (anonymous):

:-) Can be. I am not so sure about the terminology.

OpenStudy (anonymous):

with functions "inverse" means "inverse function" so the inverse of \[f(x)=x+1\] is \[f^{-1}(x)=x-1\] not \[\frac{1}{x-1}\] which is the reciprocal

OpenStudy (anonymous):

no, no graphing I just have to plug the intervals in and come up with the answer. answer in the book is, -4/pie how do I denote pie on this site?

OpenStudy (anonymous):

\[h(t)=\cot t\] itervals given, \[\pi/4, 3\pi/4\]